# mddefDefinition

### v1

@@ 3 Distribution Shifts for Machine Learning Force Fields
Figure 1: Distribution Shifts for MLFFs. We visualize distribution shifts based on changes in features, labels, and graph structure. Typical training samples from SPICE Eastman et al. (2023) and new systems from SPICEv2 (Eastman et al., 2024) are displayed. A feature shift, such as a change in elements, is shown by replacing a carbon atom with a silicon atom (left). A force norm shift is shown by the close proximity of an $H_{2}$ molecule (circled in pink), leading to high force norms (middle). A connectivity shift is shown by the tetrahedral geometry in $P_{4}S_{6}$, which differs from the typical planar geometry seen during training (right).
@@ 3.2 Criteria for Identifying Distribution Shifts
Figure 2: Distribution Shifts for Large Models. We study distribution shifts on four of the largest open-source MLFFs designed for broad chemical spaces. (a) We evaluate MACE-MP on the MPTrj train set. (b) We evaluate MACE-OFF on 10k new molecules from SPICEv2. (c) We evaluate EquiformerV2 on the OC20 out-of-distribution validation set. (d) We evaluate JMP on the ANI-1x test set. A molecule is considered out-of-distribution if it is more than 1 standard deviation away from the mean training force norm or connectivity (with respect to the spectral distance described in §3.2), or if it contains a new element. Despite their scale, these large foundation models have 210×2-10$\times$2 - 10 × larger force mean absolute errors (MAE) when encountering distribution shifts.
@@
Distribution shifts in atomic features $\mathbf{z}$ are the most apparent and detrimental to the performance of current state-of-the-art models (see §5). This may involve encountering a molecule with a new element at test time that was not present during training. For example, a model trained on $CO_{2}$ might be tested on $SiO_{2}$ without having seen $Si$ during training (see Fig. 1). Although this might initially seem like an unreasonably hard task, we argue that a truly general machine learning model for quantum chemistry should be capable of handling arbitrary elements and charges.
@@
We examine the generalization ability of MACE-OFF by testing it on 10k new molecules from the SPICEv2 dataset (Eastman et al., 2024) not included in the MACE-OFF training set. A molecule is defined as out-of-distribution if it is more than 1 standard deviation away from the mean training data force norm or connectivity (with respect to the spectral distance defined above §3.2), or if it contains a new element. Despite its scale, MACE-OFF performs worse by an order of magnitude on out-of-distribution systems (see Fig. 2a).
We evaluate JMP on the ANI-1x (Smith et al., 2020) test set defined in Shoghi et al. (2023). Although this test set does not have new elements, JMP also suffers predictably from force norm and connectivity distribution shifts (see Fig. 2d).
@@ 4.1 Test-Time Radius Refinement
Our experiments show that this procedure virtually never deteriorates performance, as one can always revert to the same radius cutoff used during training (see §5). This refinement method addresses the source of connectivity distribution shifts and serves as an efficient and effective strategy for handling new connectivities.
@@ 4.2 Test-Time Training using Cheap Priors
We further hypothesize that the current supervised training procedure for MLFFs can lead to overfitting, leading to poor representations for out-of-distribution systems and jagged potential energy landscape predictions (see Fig. 4(b) for an example on salicylic acid). To address this, we propose introducing inductive biases through improved training and inference strategies to smooth the predicted energy surfaces. The smoother energy landscape from the improved training indicates that the model may have learned more robust representations, mitigating force norm, element, and connectivity distribution shifts.
@@ 5.1 Distribution Shifts: Training on SPICE and testing on SPICEv2
Figure 6: Evaluating Distribution Shifts for Force Norms on SPICEv2. The MACE-OFF model is trained on 951k samples from the SPICE dataset, with the training force norm distribution shown in gray. We evaluate MACE-OFF on new molecules from the SPICEv2 dataset with varying force norms. As the force norms deviate further from the training distribution, MACE-OFF’s force errors increase. We also train a GemNet-T, and then apply test-time training (TTT), mitigating this shift. We highlight the top 10% of molecules with the greatest improvement to demonstrate that TTT is effective even for structures that are near the training distribution in (b).
We investigate distribution shifts from the SPICE dataset to the SPICEv2 dataset (Eastman et al., 2023, 2024) by analyzing the MACE-OFF foundation model (Kovács et al., 2023). As shown in Fig. 6, Fig. 7, and Fig. 11, we observe that despite being trained on 951k data points and scaled to 4.7M parameters, MACE-OFF experiences force norm, connectivity, and element distribution shifts when evaluated on 10k new molecules from SPICEv2 (Eastman et al., 2024). Any deviation from the training distribution, shown in gray, results in an increase in force error.
We evaluate the effectiveness of our test-time refinement strategies in mitigating these distribution shifts. For the MACE-OFF model, we implement test-time radius refinement (RR) by searching over 10 different radius cutoffs and selecting the one that best matches the training Laplacian eigenvalue distribution (see §4.1). We also train a GemNet-T model on the same training data used by MACE-OFF, using the pre-training, freezing and fine-tuning method described in §4.2, with the sGDML model as the prior (Chmiela et al., 2019). To show that TTT is prior agnostic, we additionally train a model that uses the semi-empirical GFN2-xT as the prior (Bannwarth et al., 2019). See D for more details.
@@
Both MACE-OFF and GemNet-T deteriorate in performance when encountering systems with force norms different from those seen during training, as shown in Fig. 6. Interestingly, this performance drop occurs for both higher and lower force norms than those in the training set. Test-time training reduces errors for GemNet-T on out-of-distribution force norms, and also helps decrease errors for the new systems that are closer to the training distribution. The results in Fig. 6 specifically filter out new elements and different connectivity to isolate the effect of force norm distribution shifts.
@@
For both MACE-OFF and GemNet-T, force errors increase when the connectivity of a test graph differs from that of the training graphs, as measured by the spectral distance (see Eq. 5). Our test-time radius refinement (RR) technique (see §4.1) applied to MACE-OFF effectively mitigates connectivity errors at minimal computational cost. Test-time training also effectively mitigates connectivity distribution shifts, as shown in (Fig. 7 and Tab. 4). Note that Fig. 7 isolates connectivity distribution shifts by filtering out new elements and out-of-distribution force norms. See §C.3 for RR results with the JMP model on the ANI-1x dataset.
Figure 7: Evaluating Connectivity Distribution Shifts on SPICEv2. We evaluate MACE-OFF on new molecules from the SPICEv2 dataset with varying connectivity, defined by the spectral distance to the average training graph (see §4.1 for details). Test structures with different connectivity incur larger errors for MACE-OFF. Test-time training (TTT) applied to a GemNet-T model and test-time radius refinement (RR) applied to MACE-OFF are both able to mitigate this performance drop at minimal computational cost. We highlight the top 10% of molecules with the greatest improvement to demonstrate that TTT is effective even for connectivities close to the training distribution in (b).
Elemental Distribution Shifts.
Unsurprisingly, MACE-OFF and GemNet-T perform poorly when they encounter new elements at test time. Fig. 11 shows that test-time training can reduce errors on systems with new elements, sometimes by a factor of 10 for specific molecules. While this is a challenging generalization task, we argue that achieving this should be a goal for a true chemistry foundation model, akin to first-principles methods that model the entire periodic table. Collecting more data for new elements is an option but can be prohibitively expensive, especially for atoms with many electrons. TTT provides a better starting point and reduces the amount of data that needs to be collected (see Fig. 13).
@@
We present aggregated results on the SPICEv2 distribution shift benchmark, where a model is trained on SPICE and evaluated on 10k new molecules from SPICEv2. The large MACE-OFF foundation model trains on 951k samples but still suffers from distribution shifts on the new structures from SPICEv2. We also see that (1) the RR method mitigates connectivity distribution shifts for MACE-OFF at minimal computational cost (see Tab. 1) and (2) using TTT with the GemNet-T model performs the best on the new molecules from SPICEv2, highlighting the effectiveness of training strategies for mitigating distribution shifts.
@@
SPICEv2 Test Set Force MAE (meV / Å)
With New elements No New Elements
MACE-OFF $71.2\pm 1.3$ $26.75\pm 0.65$
+RR (ours) $68.1\pm 1.6$ $26.0\pm 0.64$
GemNet-T $64.0\pm 2.5$ $22.9\pm 1.4$
+TTT (ours) $\mathbf{51.0\pm 1.8}$ $\mathbf{19.9\pm 1.0}$
Table 1: Aggregated Results on SPICEv2 Distribution Shift Benchmark. We provide aggregated results on the SPICEv2 distribution shift benchmark with 95% confidence intervals. TTT and RR are both able to effectively mitigate errors across the $10$k unseen molecules from SPICEv2.
@@ 5.2 Evaluating Generalization with Extreme Distribution Shifts: Simulating Unseen Molecules
We evaluate the stability of simulations over time by measuring deviations in bond length, following Fu et al. (2023). We additionally calculate the distribution of interatomic distances $h(r)$ to measure the quality of the simulations. See §D for more details.
@@ 6 Conclusion
We have demonstrated that state-of-the-art MLFFs, even when trained on large datasets, suffer from predictable performance degradation due to distribution shifts. By identifying shifts in element types, force norms, and connectivity, we have developed methods to diagnose the failure modes of MLFFs. Our test-time refinement methods represent initial steps in mitigating these distribution shifts, showing promising results in modeling and simulating systems outside of the training distribution. These results provide insights into how MLFFs overfit, suggesting that while MLFFs are becoming expressive enough to model diverse chemical spaces, they are not being effectively trained to do so. This may indicate that training strategies, alongside data and architecture innovations, will be important in improving MLFFs. Additionally, we have established benchmarks for evaluating the generalization ability of the next generation of MLFFs.
@@ A.2 Test-Time Radius Refinement (RR)
In this section we discuss further details about our RR approach (for theoretical justification, see §B). Although one potential worry about using RR is that it might introduce potential discontinuities, we emphasize that any model that uses a discrete set of neighbors for message passing will also experience the same issues (since a new edge could appear during simulation as soon as an atom enters the neighborhood). We note that recent work has also used k-nearest neighbor (kNN) graphs instead of (or in conjunction with) a radius cutoff (Qu & Krishnapriyan, 2024; Liao et al., 2024). However, a kNN graph can also lead to potential discontinuities from discrete neighbor changes, unless implementations explicitly account for these discontinuities (Coifman & Lafon, 2006). While it is important to continue investigating how to smooth the predicted potential energy surfaces of GNNs (Mueller et al., 2020), we emphasize that this is a problem inherent to the use of GNNs, and is not unique to the specific method of RR. Additionally, one might worry that the introduction of new edges will cause the model to overcount certain interactions. However, since edge features contain distance information, and since the model is trained on structures with varied edge distances, a well-trained model should be able to extract features from different edges. We note again that this is not an issue inherent to RR, since GNN-based MLFFs already deal with atoms entering a neighborhood during the course of simulation. Empirically, our experiments show that RR decreases force errors and improves simulation stability (see §6 and Tab. 4).
@@ Appendix C Additional Test-Time Refinement Results
Figure 11: Assessing the Impact of New Elements on Model Performance on SPICEv2 Benchmark We evaluate models trained on 951k samples from SPICE on molecules with new elements from SPICEv2. The MACE-OFF model deteriorates in performance when encountering new elements in the SPICEv2 dataset. We train a GemNet-T model on the 951k samples and run TTT—this is able to mitigate this distribution shift. We highlight the top 10% of molecules with the greatest improvement, showing that TTT can help with the challenging problem of generalizing to new elements.
@@ C.1 Further Results on SPICEv2 Distribution Shift Benchmark
We explore using the semi-empirical GFN2-xT (Bannwarth et al., 2019) as the prior to provide further evidence that TTT is agnostic of the prior chosen. We train a GemNet-dT model with the pre-train, freeze, fine-tune approach described in §4.2 using GFN2-xT as the prior. The results in Tab. 5 show that TTT with GFN2-xTB also enables better performance across a range of errors.

### v2

@@ 3 Distribution Shifts for Machine Learning Force Fields
Figure 1: Distribution Shifts for MLFFs. We visualize distribution shifts based on changes in features, labels, and graph structure. Typical training samples from SPICE Eastman et al. (2023) and new systems from SPICEv2 (Eastman et al., 2024) are displayed. An atomic feature shift is illustrated by comparing a three-atom molecule with a larger molecular system containing 91 atoms (left). A force norm shift is shown by the close proximity of an $H_{2}$ molecule (circled in pink), leading to high force norms (middle). A connectivity shift is shown by the tetrahedral geometry in $P_{4}S_{6}$, which differs from the typical planar geometry seen during training (right).
@@ 3.2 Criteria for Identifying Distribution Shifts
Figure 2: Distribution Shifts for Large Models. We study distribution shifts on four of the largest open-source MLFFs designed for broad chemical spaces. (a) We evaluate MACE-MP on the MPTrj train set. (b) We evaluate MACE-OFF on 10k new molecules from SPICEv2. (c) We evaluate EquiformerV2 on the OC20 out-of-distribution validation set. (d) We evaluate JMP on the ANI-1x test set. A molecule is considered out-of-distribution if it is more than 1 standard deviation away from the mean training force norm, system size, or connectivity (with respect to the spectral distance defined above §3.2). Despite their scale, these large foundation models have 210×2-10$\times$2 - 10 × larger force mean absolute errors (MAE) when encountering distribution shifts.
@@
Distribution shifts in atomic features refer to any change in the atomic composition of a chemical system. This includes, but is not limited to, cases where models are trained on systems containing mixtures of organic elements but tested on structures composed solely of carbon, or cases where there is a shift in system size between training and testing. Fig. 1 illustrates an atomic feature distribution shift by comparing a small carbon dioxide molecule to larger molecular system containing 91 atoms.
@@
We examine the generalization ability of MACE-OFF by testing it on 10k new molecules from the SPICEv2 dataset (Eastman et al., 2024) not included in the MACE-OFF training set. A molecule is defined as out-of-distribution if it is more than 1 standard deviation away from the mean training data force norm, system size, or connectivity (with respect to the spectral distance defined above §3.2). Despite its scale, MACE-OFF performs worse by an order of magnitude on out-of-distribution systems (see Fig. 2a). We also evaluate JMP on the ANI-1x (Smith et al., 2020) test set defined in Shoghi et al. (2023). JMP also suffers predictably from force norm, connectivity, and atomic feature distribution shifts (see Fig. 2d).
@@ 4.1 Test-Time Radius Refinement
Our experiments show that this procedure virtually never deteriorates performance, as one can always revert to the same radius cutoff used during training (see §5). Additionally, we emphasize that we update the envelope function with the new radius to ensure smoothness of the potential energy surface. We verify that this does maintain energy conservation in Tab. 3. This refinement method addresses the source of connectivity distribution shifts and serves as an efficient and effective strategy for handling new connectivities.
@@ 4.2 Test-Time Training using Cheap Priors
We further hypothesize that the current supervised training procedure for MLFFs can lead to overfitting, leading to poor representations for out-of-distribution systems and jagged potential energy landscape predictions (see Fig. 4(b) for an example on salicylic acid). To address this, we propose introducing inductive biases through improved training and inference strategies to smooth the predicted energy surfaces. The smoother energy landscape from the improved training indicates that the model may have learned more robust representations, mitigating force norm, atomic feature, and connectivity distribution shifts.
@@ 5.1 Distribution Shifts: Training on SPICE and testing on SPICEv2
Figure 5: Test-Time Training Decreases the Amount of Fine-Tuning Data Needed to Match In-Distribution Performance. We fine-tune GemNet-T models, trained on SPICE, on new molecules from the SPICEv2 dataset. Applying TTT on the new data before fine-tuning decreases the amount of training data needed to match the in-distribution performance by 10×10$\times$10 ×. Applying TTT before fine-tuning also decreases the final error by 25% when training on all the data.
We investigate distribution shifts from the SPICE dataset to the SPICEv2 dataset (Eastman et al., 2023, 2024) by analyzing the MACE-OFF foundation model (Kovács et al., 2023). As shown in Fig. 6, Fig. 7, and Fig. 11, we observe that despite being trained on 951k data points and scaled to 4.7M parameters, MACE-OFF experiences force norm, connectivity, and atomic feature distribution shifts when evaluated on 10k new molecules from SPICEv2 (Eastman et al., 2024). Any deviation from the training distribution, shown in gray, predictably results in an increase in force error.
We evaluate the effectiveness of our test-time refinement strategies in mitigating these distribution shifts. For the MACE-OFF model, we implement test-time radius refinement (RR) by searching over 10 different radius cutoffs and selecting the one that best matches the training Laplacian eigenvalue distribution (see §4.1). We also train a GemNet-T model on the same training data used by MACE-OFF, using the pre-training, freezing and fine-tuning method described in §4.2, with the sGDML model as the prior (Chmiela et al., 2019). To show that TTT is prior agnostic, we additionally train a model that uses the semi-empirical GFN2-xTB as the prior (Bannwarth et al., 2019). See D for more details.
@@
Both MACE-OFF and GemNet-T deteriorate in performance when encountering systems with force norms different from those seen during training, as shown in Fig. 6. Interestingly, this performance drop occurs for both higher and lower force norms than those in the training set. Test-time training reduces errors for GemNet-T on out-of-distribution force norms, and also helps decrease errors for the new systems that are closer to the training distribution. The results in Fig. 6 specifically filter out atomic feature shifts and different connectivities to isolate the effect of force norm distribution shifts.
@@
For both MACE-OFF and GemNet-T, force errors increase when the connectivity of a test graph differs from that of the training graphs, as measured by the spectral distance (see Eq. 5). Our test-time radius refinement (RR) technique (see §4.1) applied to MACE-OFF effectively mitigates connectivity errors at minimal computational cost. Test-time training also effectively mitigates connectivity distribution shifts, as shown in (Fig. 7 and Tab. 5). Note that Fig. 7 isolates connectivity distribution shifts by filtering out-of-distribution force norms and atomic features. See §C.3 for RR results with the JMP model on the ANI-1x dataset.
Figure 6: Evaluating Distribution Shifts for Force Norms on SPICEv2. We evaluate MACE-OFF on new molecules from the SPICEv2 dataset with varying force norms. (a) Test structures with different force norms relative to the training distribution (shown in gray) incur larger force errors for MACE-OFF. We also train a GemNet-T model, and then apply test-time training (TTT), mitigating this shift. (b) We highlight the top 10% of molecules with the greatest improvement to demonstrate that TTT is effective even for structures that are near the training distribution.
Figure 7: Evaluating Connectivity Distribution Shifts on SPICEv2. We evaluate MACE-OFF on new molecules from the SPICEv2 dataset with varying connectivity, defined by the spectral distance to the average training graph (see §4.1 for details). (a) Test structures with different connectivity relative to the training distribution (shown in gray) incur larger force errors for MACE-OFF. Test-time training (TTT) applied to a GemNet-T model and test-time radius refinement (RR) applied to MACE-OFF are both able to mitigate this performance drop at minimal computational cost. (b) We highlight the top 10% of molecules with the greatest improvement to demonstrate that TTT is effective even for connectivities close to the training distribution.
Atomic Feature Distribution Shifts.
MACE-OFF and GemNet-T both perform poorly when encountering molecules with atomic features that differ from their training distributions. In particular, MACE-OFF and GemNet-T struggle with molecules that are both larger and smaller than those seen in training, and with systems that have a different proportion of carbon atoms than seen in training (see Fig. 11). Test-time training reduces errors across both of these atomic feature distribution shifts for GemNet-T. We filter out out-of-distribution connectivities and force norms to isolate the effect of atomic feature distribution shifts.
@@
We present aggregated results on the SPICEv2 distribution shift benchmark, where a model is trained on SPICE and evaluated on 10k new molecules from SPICEv2. The large MACE-OFF foundation model trains on 951k samples but still suffers from distribution shifts on the new structures from SPICEv2. We also see that (1) the RR method mitigates connectivity distribution shifts for MACE-OFF at minimal computational cost (see Tab. 1) and (2) using TTT with the GemNet-T model performs the best on the new molecules from SPICEv2, highlighting the effectiveness of training strategies for mitigating distribution shifts. Practically, these lower force errors also translate into better MD simulations and improved structure relaxations (see §C.1).
@@
Model SPICEv2 Test Set Force MAE (meV/Å)
MACE-OFF $26.75\pm 0.65$
+RR (ours) $26.0\pm 0.64$
GemNet-T $22.9\pm 1.4$
+TTT (ours) $\mathbf{19.9\pm 1.0}$
Table 1: Aggregated Results on SPICEv2 Distribution Shift Benchmark. We provide aggregated results on the SPICEv2 distribution shift benchmark with 95% confidence intervals. TTT and RR are both able to effectively mitigate errors across the $10$k unseen molecules from SPICEv2. The relative improvements observed are in line with previous test-time training work (Gandelsman et al., 2022; Sun et al., 2020).
@@
Test-Time Training and Fine-Tuning.
While TTT can enable accurate MD simulations for new systems without access to any reference labels (see §5.2 and §C.1), a practitioner may prefer to fine-tune a model on more out-of-distribution data to match the in-distribution performance. We examine how TTT can provide a better starting point for fine-tuning by learning more robust representations for new systems.
We take the GemNet-T models from the previous section and fine-tune them with varying amounts of structures from the SPICEv2 dataset (Eastman et al., 2024). We evaluate how much data is required to match the in-distribution force error on SPICE ($12.9$ meV/Å) when tested on the 10k new molecules from SPICEv2. The vanilla GemNet-T model matches the in-distribution performance when trained on half of the SPICEv2 data. In contrast, using our TTT procedure before fine-tuning allows the model to reach the same performance with only $5\%$ of the data — a 10×10$\times$10 × reduction. Additionally, TTT reduces the final error by $25\%$ even when fine-tuning on the entire SPICEv2 dataset. See Fig. 5 for details.
@@ 5.2 Evaluating Generalization with Extreme Distribution Shifts: Simulating Unseen Molecules
We evaluate the stability of simulations over time by measuring deviations in bond length, following Fu et al. (2023). We additionally calculate the distribution of interatomic distances $h(r)$, a low dimensional descriptor of 3D molecular structures, to measure the quality of the simulations (Zhang et al., 2018; Fu et al., 2023; Raja et al., 2025). See §D for more details.
@@ 6 Conclusion
We have demonstrated that state-of-the-art MLFFs, even when trained on large datasets, suffer from predictable performance degradation due to distribution shifts. By identifying shifts in atomic features, force norms, and connectivity, we have developed methods to diagnose the failure modes of MLFFs. Our test-time refinement methods represent initial steps in mitigating these distribution shifts, showing promising results in modeling and simulating systems outside of the training distribution. These results provide insights into how MLFFs overfit, suggesting that while MLFFs are becoming expressive enough to model diverse chemical spaces, they are not being effectively trained to do so. This may indicate that training strategies, alongside data and architecture innovations, will be important in improving MLFFs. Finally, our experiments serve as benchmarks for evaluating the generalization ability of the next generation of MLFFs.
@@ A.2 Test-Time Radius Refinement (RR)
In this section we discuss further details about our RR approach (for theoretical justification, see §B). Although one potential worry about using RR is that it might introduce potential discontinuities, we emphasize that we update the envelope function to ensure that the predicted potential energy surface remains smooth with the new radius. When running MD simulations, we choose the updated radius at the beginning and keep it fixed over the course of simulation. We verify that this maintains a conservative force field by running NVE simulations (see Tab. 3). Additionally, one might worry that the introduction of new edges will cause the model to overcount certain interactions. However, since edge features contain distance information, and since the model is trained on structures with varied edge distances, a well-trained model should be able to extract features from different edges. We note again that this is not an issue inherent to RR, since GNN-based MLFFs already deal with atoms entering a neighborhood during the course of simulation. Empirically, our experiments show that RR decreases force errors and improves simulation stability (see §5 and Tab. 5).
Model Energy Deviation (eV)
MACE-OFF 0.0036 ± 0.0004
MACE-OFF + RR 0.0049 ± 0.0022
GemNet-dT (non-conservative) >1.0
Table 3: RR Maintains Energy Conservation. When using RR, we select the updated radius for the new molecule at the start of simulation and then keep it fixed. We update the envelope function to ensure smoothness of the predicted potential energy surface with the new radius. We run 10 ps NVE simulations to verify that RR does maintain a conservative force field.
@@ Appendix C Additional Test-Time Refinement Results
Figure 11: Assessing the Impact of Atomic Feature Distribution Shifts on Model Performance on SPICEv2 Benchmark We evaluate models trained on 951k samples from SPICE on new molecules from SPICEv2. The MACE-OFF model deteriorates in performance when encountering systems of different sizes or molecules with a different proportion of carbons atoms compared to its training set. We train a GemNet-T model on the 951k samples and run TTT—this is able to mitigate the atomic feature distribution shifts.
@@ C.1 Further Results on SPICEv2 Distribution Shift Benchmark
Evaluating Improved Downstream Utility.
We run reference DFT simulations to ensure that the improved stability reported in Tab. 4 and Tab. 5 translate into improved simulation quality in terms of the predicted distribution of interatomic distances h(r) (see §5.2 and §D for details). Due to the computational cost of running reference MD simulations, we are only able to do this on a subset of the molecules. RR on top of MACE-OFF lowers the h(r) MAE from $0.17$ and $0.09$ to $0.15$ and $0.07$ for $SC_{6}H_{4}$ and $IC_{2}H$, respectively. TTT on top of GemNet-T lowers the h(r) MAE from $0.20,$ $0.44,$ and $0.39$ to $0.18,$ $0.19,$ and $0.17$ for $N_{3}C_{5}H_{3}$, $IC_{2}H$, and $O_{3}P$, respectively.
We run also BFGS structure relaxations with the MLFFs and caclulate how many extra steps are needed with DFT to find a relaxed structure. GemNet-T requires 10.4 additional steps on average (on $N_{3}C_{5}H_{3}$, $IC_{2}H$, and $O_{3}P$, 3 seeds each), whereas GemNet-T+TTT requires 7.7. MACE-OFF requires 4.0, and MACE-OFF+RR requires 3.8 (on $SC_{6}H_{4}$ and $IC_{2}H$, 3 seeds each).
@@
We explore using the semi-empirical GFN2-xTB (Bannwarth et al., 2019) as the prior to provide further evidence that TTT is agnostic of the prior chosen. We train a GemNet-dT model with the pre-train, freeze, fine-tune approach described in §4.2 using GFN2-xTB as the prior. The results in Tab. 6 show that TTT with GFN2-xTB also enables better performance across a range of errors.